Which equation represents a line which is perpendicular to the line y=-4x-5?
x+4y=−32
4x-y=-4
4x+y=3
4y-x=4

Answer by mananth(16946)

y=-4x-5
sLOPE =-4


4y-x=4
4y=x+4
y=1/4 x +1
4y-x=4 is perpendicular